Part A: The change in enthalpy (ΔHorxn) for a reaction is -25.7 kJ/mol . The equilibrium constant for the reaction is 1.3×103 at 298 K. What is the equilibrium constant for the reaction at 668 K?

Part B: A reaction has an equilibrium constant of 7.4×103 at 298 K. At 753 K , the equilibrium constant is 0.63. Find ΔHorxn for the reaction.
Answer #1

A)

Given:

T1 = 298 K

T2 = 668 K

K1 = 1.3*10^3

Ho = -25.7 KJ/mol

= -25700 J/mol

use:

ln(K2/K1) = (Ho/R)*(1/T1 - 1/T2)

ln(K2/1.3*10^3) = (-25700.0/8.314)*(1/298 - 1/668.0)

ln(K2/1.3*10^3) = -3091*(1.859*10^-3)

ln(K2/1.3*10^3) = -5.746

(K2/1.3*10^3) = e^(-5.746)

(K2/1.3*10^3) = 3.197*10^-3

K2 = 4.156

Answer: 4.16

B)

Given:

T1 = 298 K

T2 = 753 K

K1 = 7.4*10^3

K2 = 0.63

use:

ln(K2/K1) = (Ho/R)*(1/T1 - 1/T2)

ln(0.63/7.4*10^3) = ( Ho/8.314)*(1/298 - 1/753)

-9.3713 = (Ho/8.314)*(2.028*10^-3)

Ho = -38425 J/mol

Ho = -38.4245 KJ/mol

Answer: -38.4 KJ/mol